OpenStudy (anonymous):

How do I write a system of equations with the solution (4,-3)?

OpenStudy (anonymous):

First, write any system of 2 linear equations you like on the left side. For example: 3x - 2y = 4x + 3y = Now, find the right side by replacing x = 4 and y = -3 in the 2 equations you wrote. Watch out the sign of -y. 3(4) - 2(-3) = 12 + 6 = ..... 4(4) + 3(-3) = .......